How many liters of whole milk containing 3.5% butterfat must be mixed with 3 liters of skim milk with 0% butterfat to obtain a mixture containing 2% butterfat?
How would I write this in an equation?
My guess is 3 + .035x = .02(3+x)

Answer by stanbon(75887)

How many liters of whole milk containing 3.5% butterfat must be mixed with 3 liters of skim milk with 0% butterfat to obtain a mixture containing 2% butterfat?
How would I write this in an equation?
My guess is 3 + .035x = .02(3+x)
----------------------------------
Equation:
fat + fat = fat
0*3 + 0.035x = 0.02(3+x)
---
Multiply thru by 1000 to get:
0 + 35x = 20*3 + 20x
15x = 20*3
x = 4 liters (amt of 3.5% milk needed in the mixture)
=============================
Cheers,
Stan H.